# Future targets for light gauge bosons from cosmic strings

### v1

@@ Abstract
I Introduction
Cosmic strings are hypothetical one-dimensional topological defects believed to have formed during the very early stages of the Universe, potentially as a result of primordial phase transitions Kibble (1976); Vilenkin (1981, 1985); Witten (1985). Despite being purely theoretical, cosmic strings are of great interest in cosmology due to their potential influence on the large-scale structure of the Universe and their possible observable effects. For instance, cosmic strings could seed the formation of galaxies and clusters in the early Universe, potentially leaving distinct imprints on the cosmic microwave background and large-scale matter distribution that would provide clues to their existence Brandenberger (1991, 1994); Charnock et al. (2016); Fernandez et al. (2020). Cosmic strings could also explain the baryon asymmetry through leptogenesis Bhattacharjee et al. (1982) and trigger the formation of primordial black holes Hawking (1989); Fort and Vachaspati (1993).
@@ I Introduction
The evolution of the network of cosmic strings related with the spontaneous breaking of a global symmetry has long been discussed since its early establishment Vilenkin (1985); Witten (1985); Vachaspati and Vilenkin (1985); Davis (1985, 1986); Vilenkin and Vachaspati (1987); Harari and Sikivie (1987); Davis and Shellard (1989); Dabholkar and Quashnock (1990); Hagmann and Sikivie (1991); Allen and Shellard (1992); Battye and Shellard (1994); Davis and Peter (1995). A key property of such a global string network is the development of a scaling regime in which each Hubble volume contains around one long string Hindmarsh and Kibble (1995); Vilenkin and Shellard (2000). The scaling is assured by the continuous production of small string loops through the chopping and reconnection of long strings.
@@
Recent advancements in global string simulations have enabled the study of gravitational wave (GW) emissions from a network of global strings Gorghetto et al. (2018, 2021); Chang and Cui (2022). A significant characteristic of the GW background (GWB) produced by these strings is its emission across a wide range of frequencies as the string network evolves. This distinctive feature sets the GWB from strings apart from other possible sources, such as cosmological phase transitions or inflation, which may also have been active before Big Bang Nucleosynthesis (BBN) Caprini and Figueroa (2018); Giovannini (1998); Bernal and Hajkarim (2019); D’Eramo and Schmitz (2019); Vagnozzi and Loeb (2022); Caprini et al. (2024). For this, the GWB predicted from new physics has been vigorously studied to pin down the changes brought in by a modified pre-BBN thermal history, and it is actively sought in experiments Cui et al. (2018, 2019); Guedes et al. (2018); Gouttenoire et al. (2020); Auclair et al. (2020); Ramberg and Visinelli (2019, 2021); Buchmuller et al. (2021); Ghoshal et al. (2023); Servant and Simakachorn (2024); Baeza-Ballesteros et al. (2024a); Schmitz and Schröder (2024). One important example of a global string network is related to the breaking of the global U(1) Peccei-Quinn symmetry invoked to solve the strong CP problem Peccei and Quinn (1977a, b) and possibly leading to the emission of pseudo-Nambu Goldstone bosons from cosmic strings Davis (1986); Vilenkin and Vachaspati (1987); Harari and Sikivie (1987); Davis and Shellard (1989); Dabholkar and Quashnock (1990); Hagmann and Sikivie (1991); Allen and Shellard (1992); Battye and Shellard (1994).
@@
One alternative possibility involves the breaking of a gauge symmetry, leading to the formation of a cosmic string network composed of gauge strings. A gauge string solution shares similarities with a global string, although the energy from the gradient term outside the core regime exhibits finite tension concentrated within the string core. While loops with lengths below a critical value predominantly decay through particle emission, larger loops shrink via the release of gravitational waves (GWs), similar to the global case Matsunami et al. (2019); Baeza-Ballesteros et al. (2024b).
@@
The production of gauge bosons from the decay of cosmic strings has gained considerable interest in recent years, as it represents a successful model to produce such particles as dark matter in the early Universe Jeannerot et al. (1999); Lin et al. (2001); Matsuda (2006); Cui and Morrissey (2009); Kuroyanagi et al. (2013); Long and Wang (2019); Kitajima and Nakayama (2023a, b). Alternative mechanisms may suffer some issues. For example, the production from misalignment requires some degree of fine tuning unless the gauge boson couples non-minimally to gravity Arias et al. (2012); Graham et al. (2016); Alonso-Álvarez et al. (2020); Caputo et al. (2021). Relativistic gauge bosons radiated from strings are also expected to contribute to the dark radiation budget to some degree, in particular this component can be constrained by the considerations over the number of relativistic neutrinos that affect free-streaming at recombination. Other models that have been recently considered involve metastable cosmic strings Kibble (1976); Vilenkin (1982); Antusch et al. (2024).
@@
In this work, we explore the simultaneous production of gauge bosons and the GWB by a local string network. We explicitly consider the Abelian Higgs model, where the U(1) symmetry predicts the existence of a massive gauge field that could contribute as a cold component of the Universe. We compare analytical approaches from the literature and derive the frequency dependence of the GWB for different regimes. These analytical derivations are useful for numerical tests of the string network. While these results have been somewhat discussed in previous literature, we have delved deeper into the details of the expressions for the GWB and compiled the derivations of GW abundances across various frequency ranges. For completeness, we also compare our results with the sensitivities of planned GW detectors. We work with units $\hbar=c=1$.
II Methods
II.1 The Abelian Higgs Model
@@
The Lagrangian in Eq. (2) allows for a solution in terms of a cosmic string, a linear concentration of energy of cosmological size which is produced along with a cosmological phase transition in the early Universe Nielsen and Olesen (1973), see also Refs. Hindmarsh and Kibble (1995); Vilenkin and Shellard (2000). The string solution is described below.
II.2 Gravitational wave emission from cosmic strings
The equation of motion associated with the theory in Eq. (2) allows for a solution describing a cosmic string Vilenkin (1981, 1985). The string configuration is described by a scalar field displaced from its minimum within a narrow region of radius $r\leq 1/m_{\rho}$ and decaying in the outer region to the horizon scale $r\sim 1/H(t)$. The non-zero spatial gradient in the outer region holds most of the energy density of the field, so that the string tension that arises from the spatial gradient of the scalar field’s phase depends logarithmically on the ratio of the ultraviolet (UV) scale $1/m_{\rho}$ and the infrared (IR) scale $\sim 1/H$ as
@@ II.2 Gravitational wave emission from cosmic strings
The set of cosmic strings develop into a network at energies below the spontaneous symmetry breaking scale $v$, with strings showing a complex dynamics through mutual intersections and reconnections. Within a Hubble time, cosmic strings intersections form new structures such as loops, kinks and cusps Vachaspati and Vilenkin (1985); Garfinkle and Vachaspati (1987); Damour and Vilenkin (2000, 2001); Olmez et al. (2010) which dissolve by the emission of strong bursts of gravitational waves (GWs).
@@
where $\Gamma_{\rm GW}$ is a dimensionless constant that describes the share of energy that a single smooth closed loop releases in GWs and the Newton’s constant is $G=1/m_{\rm Pl}^{2}$ in terms of Plank’s mass $m_{\rm Pl}$. Numerical simulations fix this constant as $\Gamma_{\rm GW}\simeq 50$ Blanco-Pillado et al. (2014). Oscillating closed string loops are typically the dominant source of GWs from a cosmic string network in the scaling regime. For a string of energy $E_{\ell}=\mu\ell$ and neglecting the time dependence of the string tension in the integration, the length $\ell=\ell(t)$ of a string loop created by the network at time $t_{i}$ evolves according to
@@
The stochastic GWB depends on the rate of loop production by the cosmic string network. We model this using the velocity-dependent one-scale (VOS) model Martins and Shellard (1996a, b, 2002). The energy of the network transferred to the loops is defined as
@@
where $C_{\rm loop}$ parameterizes the fraction of energy transferred from the long string network into large loops and $\rho_{\rm str}$ is the string energy density. Here, we fix $C_{\rm loop}\approx 10\%$ according to recent cosmic string simulations Blanco-Pillado et al. (2014); Blanco-Pillado and Olum (2017), which implies that the remaining 90% of the energy is transferred to the kinetic energies of highly-boosted smaller loops. Applying local energy conservation and considering only large loops, this yields a loop formation rate per unit volume at time $t_{i}$ of
@@
The function $C_{\rm eff}$ depends on the redshift scaling of the dominant energy density $\rho$ of the universe and it is obtained within the VOS model as $C_{\rm eff}^{\rm MD}=0.41$ for matter domination (MD) and $C_{\rm eff}^{\rm RD}=5.5$ for radiation domination (RD) Cui et al. (2018). Here, we interpolate between these results by adopting the expression
@@
where the last factor accounts for the redshift of radiation between emission and today, and the average power (energy per unit time) of GWs emitted by a single loop with mode $k$ is modeled in terms of a function $\Gamma_{k}$ that accounts for the emission details as Vilenkin (1981); Turok (1984); Vachaspati and Vilenkin (1985); Burden (1985); Garfinkle and Vachaspati (1987)
@@
We consider the GW emission from cusps, which is described by Vachaspati and Vilenkin (1985); Garfinkle and Vachaspati (1987); Damour and Vilenkin (2000, 2001); Olmez et al. (2010)
@@
where $\rho_{c}=3H_{0}^{2}/(8\pi G)$ is the present critical density. The GW energy density per unit logarithmic frequency emitted from the formation of the string network at time $t_{F}\sim 1/(2v)$ to the present time is given by Cui et al. (2008, 2018, 2019)
@@
where $t_{i,{\rm GW}}$ is the formation time obtained from Eq. (9) and $\Theta$ is the Heaviside theta function. When the string network emits only GWs, mimicking a global string network, the GW abundance in Eq. (18) at high frequencies predicts $\Omega_{\rm GW}(f)\propto\sqrt{\mu}$ Blanco-Pillado et al. (2014).
II.3 The role of gauge bosons emission
@@
where the dimensionless constant $\Gamma_{\rm vec}\sim 65$ parameterizes the strength of the release in gauge bosons, with this numerical value determined by numerical simulations Blanco-Pillado et al. (2014). The step function in Eq. (6) ensures that the loop cannot emit gauge bosons if the inverse of the initial loop size is smaller than the gauge boson mass. The length scale $1/m_{A^{\prime}}$ introduces a new cutoff scale, which could potentially be smaller than the horizon scale, thereby modifying the string tension in Eq. (5) as Gorghetto et al. (2018)
@@
II.4 Relic gauge boson abundance
The evolution of the string energy density is governed by the release of gauge boson emission at early times, according to the expression Harari and Sikivie (1987)
@@ II.4 Relic gauge boson abundance
Here, $\xi$ is the average string number per Hubble volume, which appears to show a logarithmic violation of the scaling law as $\xi=\xi_{0}+\xi_{1}\ln(m_{\rho}/H)$ in some numerical simulations Kawasaki et al. (2018); Vaquero et al. (2019); Gorghetto et al. (2018); Buschmann et al. (2020); Klaer and Moore (2020); Gorghetto et al. (2021); Buschmann et al. (2022), with other studies claiming the conventional scaling law $\xi\approx 1$ Yamaguchi et al. (1999); Yamaguchi (1999); Yamaguchi et al. (2000); Hagmann et al. (2001); Hiramatsu et al. (2011, 2012); Kawasaki et al. (2015); Lopez-Eiguren et al. (2017); Hindmarsh et al. (2020, 2021). Here, we adopt the former approach and consider the average number of long strings per Hubble volume showing a logarithmic dependence on the gauge boson mass as $\xi=\xi_{1}\ln(m_{\rho}/m_{A^{\prime}})$, where we fix $\xi_{1}=0.15$.
@@
Here, the integration runs over the magnitude of the momentum of the radiated particles $k$. In the assumption that only the modes with energy $E_{k}\sim H(t)$ contribute, the source density reads $\mathcal{S}_{A}\approx\mathcal{P}/H(t)$, so that the solution of Eq. (30) leads to the present abundance of non-relativistic gauge bosons as Kitajima and Nakayama (2023a)
@@
Note, that alternative derivations in the literature find the same power-law dependence, despite differing by an overall numerical prefactor Gorghetto et al. (2018); Long and Wang (2019).
Along with the cold component, the network emits relativistic gauge bosons that add up to the dark radiation budget in the cosmic microwave background (CMB). The contribution at times $t<t_{A^{\prime}}$ leads to Gorghetto et al. (2021)
@@
In principle, Eq. (36) should be modified to account for the effects of a non-zero mass at CMB Caloni et al. (2022). We do not consider such results here because the bound on $m_{A^{\prime}}$ resulting from Eq. (32) imposes that the gauge boson be lighter than the temperatures in play at CMB.
II.5 Cosmological setup
To derive our results below, we fix the cosmological parameters according to the best fit by the Planck collaboration for the TT,TE,EE+lowE+lensing datasets Aghanim et al. (2020). More in detail, the standard cosmological model parameters are set by taking the Hubble constant as $H_{0}=100\,h{\rm\,km\,s^{-1}\,Mpc^{-1}}$ with $h=0.674$. The redshift at matter-radiation equality is approximately $z_{\rm eq}\approx 3402$. The matter density parameter is fixed as $\Omega_{m}=0.3153$, leading to a radiation density parameter of $\Omega_{r}=\Omega_{m}/(1+z_{\rm eq})=9.3\times 10^{-5}$. The dark energy density parameter is then $\Omega_{\Lambda}\approx 0.685$. The scale factor as a function of cosmic time can be obtained by integrating the Friedmann equation

### v2

@@ Abstract
journal: Physics of the Dark Universe
1 Introduction
Cosmic strings are hypothetical one-dimensional topological defects believed to have formed during the very early stages of the Universe, potentially as a result of primordial phase transitions [1, 2, 3, 4]. Despite being purely theoretical, cosmic strings are of great interest in cosmology due to their potential influence on the large-scale structure of the Universe and their possible observable effects. For instance, cosmic strings could seed the formation of galaxies and clusters in the early Universe, potentially leaving distinct imprints on the cosmic microwave background and large-scale matter distribution that would provide clues to their existence [5, 6, 7, 8]. Cosmic strings could also explain the baryon asymmetry through leptogenesis [9] and trigger the formation of primordial black holes [10, 11].
@@ 1 Introduction
The evolution of the network of cosmic strings related with the spontaneous breaking of a global symmetry has long been discussed since its early establishment [3, 4, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. A key property of such a global string network is the development of a scaling regime in which each Hubble volume contains around one long string [23, 24]. The scaling is assured by the continuous production of small string loops through the chopping and reconnection of long strings.
@@
Recent advancements in global string simulations have enabled the study of gravitational wave (GW) emissions from a network of global strings [25, 26, 27]. A significant characteristic of the GW background (GWB) produced by these strings is its emission across a wide range of frequencies as the string network evolves. This distinctive feature sets the GWB from strings apart from other possible sources, such as cosmological phase transitions or inflation, which may also have been active before Big Bang Nucleosynthesis (BBN) [28, 29, 30, 31, 32, 33]. For this, the GWB predicted from new physics has been vigorously studied to pin down the changes brought in by a modified pre-BBN thermal history, and it is actively sought in experiments [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. One important example of a global string network is related to the breaking of the global U(1) Peccei-Quinn symmetry invoked to solve the strong CP problem [48, 49] and possibly leading to the emission of pseudo-Nambu Goldstone bosons from cosmic strings [14, 15, 16, 17, 18, 19, 20, 21].
@@
One alternative possibility involves the breaking of a gauge symmetry, leading to the formation of a cosmic string network composed of gauge strings. A gauge string solution shares similarities with a global string, although the energy from the gradient term outside the core regime exhibits finite tension concentrated within the string core. While loops with lengths below a critical value predominantly decay through particle emission, larger loops shrink via the release of gravitational waves (GWs), similar to the global case [50, 51].
@@
The production of gauge bosons from the decay of cosmic strings has gained considerable interest in recent years, as it represents a successful model to produce such particles as dark matter in the early Universe [52, 53, 54, 55, 56, 57, 58, 59]. Alternative mechanisms may suffer some issues. For example, the production from misalignment requires some degree of fine tuning unless the gauge boson couples non-minimally to gravity [60, 61, 62, 63]. Relativistic gauge bosons radiated from strings are also expected to contribute to the dark radiation budget to some degree, in particular this component can be constrained by the considerations over the number of relativistic neutrinos that affect free-streaming at recombination. Other models that have been recently considered involve metastable cosmic strings [1, 64, 65].
@@
In this work, we explore the simultaneous production of gauge bosons and the GWB by a local string network. We explicitly consider the Abelian Higgs model, where the U(1) symmetry predicts the existence of a massive gauge field that could contribute as a cold component of the Universe. We compare analytical approaches from the literature and derive the frequency dependence of the GWB for different regimes. These analytical derivations are useful for numerical tests of the string network. These results have been somewhat discussed in previous literature, see e.g. Refs. [57, 58] where the evolution of an Abelian string network is studied with numerical simulations; here, we have delved deeper into the details of the expressions for the GWB and compiled the derivations of GW abundances across various frequency ranges. For completeness, we also compare our results with the sensitivities of planned GW detectors. We work with units $\hbar=c=1$.
2 Methods
2.1 The Abelian Higgs Model
@@
The Lagrangian in Eq. (2) allows for a solution in terms of a cosmic string, a linear concentration of energy of cosmological size which is produced along with a cosmological phase transition in the early Universe [66], see also Refs. [23, 24]. The string solution is described below.
2.2 Gravitational wave emission from cosmic strings
The equation of motion associated with the theory in Eq. (2) allows for a solution describing a cosmic string [2, 3]. The string configuration is described by a scalar field displaced from its minimum within a narrow region of radius $r\leq 1/m_{\rho}$ and decaying in the outer region to the horizon scale $r\sim 1/H(t)$. The non-zero spatial gradient in the outer region holds most of the energy density of the field, so that the string tension that arises from the spatial gradient of the scalar field’s phase depends logarithmically on the ratio of the ultraviolet (UV) scale $1/m_{\rho}$ and the infrared (IR) scale $\sim 1/H$ as
@@ 2.2 Gravitational wave emission from cosmic strings
The set of cosmic strings develop into a network at energies below the spontaneous symmetry breaking scale $v$, with strings showing a complex dynamics through mutual intersections and reconnections. Within a Hubble time, cosmic strings intersections form new structures such as loops, kinks and cusps [12, 67, 68, 69, 70] which dissolve by the emission of strong bursts of gravitational waves (GWs).
@@
where $\Gamma_{\rm GW}$ is a dimensionless constant that describes the share of energy that a single smooth closed loop releases in GWs and the Newton’s constant is $G=1/m_{\rm Pl}^{2}$ in terms of Plank’s mass $m_{\rm Pl}$. Numerical simulations fix this constant as $\Gamma_{\rm GW}\simeq 50$ [71]. Oscillating closed string loops are typically the dominant source of GWs from a cosmic string network in the scaling regime. For a string of energy $E_{\ell}=\mu\ell$ and neglecting the time dependence of the string tension in the integration, the length $\ell=\ell(t)$ of a string loop created by the network at time $t_{i}$ evolves according to
@@
The stochastic GWB depends on the rate of loop production by the cosmic string network. We model this using the velocity-dependent one-scale (VOS) model [72, 73, 74]. The energy of the network transferred to the loops is defined as
@@
where $C_{\rm loop}$ parameterizes the fraction of energy transferred from the long string network into large loops and $\rho_{\rm str}$ is the string energy density. Here, we fix $C_{\rm loop}\approx 10\%$ according to recent cosmic string simulations [71, 75], which implies that the remaining 90% of the energy is transferred to the kinetic energies of highly-boosted smaller loops. Applying local energy conservation and considering only large loops, this yields a loop formation rate per unit volume at time $t_{i}$ of
@@
The function $C_{\rm eff}$ depends on the redshift scaling of the dominant energy density $\rho$ of the universe and it is obtained within the VOS model as $C_{\rm eff}^{\rm MD}=0.41$ for matter domination (MD) and $C_{\rm eff}^{\rm RD}=5.5$ for radiation domination (RD) [34]. Here, we interpolate between these results by adopting the expression
@@
where the last factor accounts for the redshift of radiation between emission and today, and the average power (energy per unit time) of GWs emitted by a single loop with mode $k$ is modeled in terms of a function $\Gamma_{k}$ that accounts for the emission details as [2, 76, 12, 77, 67]
@@
We consider the GW emission from cusps, which is described by [12, 67, 68, 69, 70]
@@
where $\rho_{c}=3H_{0}^{2}/(8\pi G)$ is the present critical density. The GW energy density per unit logarithmic frequency emitted from the formation of the string network at time $t_{F}\sim 1/(2v)$ to the present time is given by [78, 34, 35]
@@
where $t_{i,{\rm GW}}$ is the formation time obtained from Eq. (9) and $\Theta$ is the Heaviside theta function. When the string network emits only GWs, mimicking a global string network, the GW abundance in Eq. (18) at high frequencies predicts $\Omega_{\rm GW}(f)\propto\sqrt{\mu}$ [71].
2.3 The role of gauge bosons emission
@@
where the dimensionless constant $\Gamma_{\rm vec}\sim 65$ parameterizes the strength of the release in gauge bosons, with this numerical value determined by numerical simulations [71]. The step function in Eq. (6) ensures that the loop cannot emit gauge bosons if the inverse of the initial loop size is smaller than the gauge boson mass. The length scale $1/m_{A^{\prime}}$ introduces a new cutoff scale, which could potentially be smaller than the horizon scale, thereby modifying the string tension in Eq. (5) as [25]
@@
Solving Eq. (27) allows for the reconstruction of the GW spectrum as a function of frequency. In particular, when compared with numerical results (see, e.g., Ref. [59]), our method reproduces the same pivot frequencies and spectral slopes of the GW spectrum, although it generally predicts peak amplitudes that are smaller by about an order of magnitude.
2.4 Relic gauge boson abundance
The evolution of the string energy density is governed by the release of gauge boson emission at early times, according to the expression [16]
@@ 2.4 Relic gauge boson abundance
Here, $\xi$ is the average string number per Hubble volume, which appears to show a logarithmic violation of the scaling law as $\xi=\xi_{0}+\xi_{1}\ln(m_{\rho}/H)$ in some numerical simulations [79, 80, 25, 81, 82, 26, 83], with other studies claiming the conventional scaling law $\xi\approx 1$ [84, 85, 86, 87, 88, 89, 90, 91, 92, 93]. Here, we adopt the former approach and consider the average number of long strings per Hubble volume showing a logarithmic dependence on the gauge boson mass as $\xi=\xi_{1}\ln(m_{\rho}/m_{A^{\prime}})$, where we fix $\xi_{1}=0.15$.
@@
Here, the integration runs over the magnitude of the momentum of the radiated particles $k$. In the assumption that only the modes with energy $E_{k}\sim H(t)$ contribute, the source density reads $\mathcal{S}_{A}\approx\mathcal{P}/H(t)$, so that the solution of Eq. (30) leads to the present abundance of non-relativistic gauge bosons as [58]
@@
Note, that alternative derivations in the literature find the same power-law dependence, despite differing by an overall numerical prefactor [25, 57].
Along with the cold component, the network emits relativistic gauge bosons that add up to the dark radiation budget in the cosmic microwave background (CMB). The contribution at times $t<t_{A^{\prime}}$ leads to [26]
@@
In principle, Eq. (36) should be modified to account for the effects of a non-zero mass at CMB [94]. We do not consider such results here because the bound on $m_{A^{\prime}}$ resulting from Eq. (32) imposes that the gauge boson be lighter than the temperatures in play at CMB.
2.5 Cosmological setup
To derive our results below, we fix the cosmological parameters according to the best fit by the Planck collaboration for the TT,TE,EE+lowE+lensing datasets [95]. More in detail, the standard cosmological model parameters are set by taking the Hubble constant as $H_{0}=100\,h{\rm\,km\,s^{-1}\,Mpc^{-1}}$ with $h=0.674$. The redshift at matter-radiation equality is approximately $z_{\rm eq}\approx 3402$. The matter density parameter is fixed as $\Omega_{m}=0.3153$, leading to a radiation density parameter of $\Omega_{r}=\Omega_{m}/(1+z_{\rm eq})=9.3\times 10^{-5}$. The dark energy density parameter is then $\Omega_{\Lambda}\approx 0.685$. The scale factor as a function of cosmic time can be obtained by integrating the Friedmann equation